# Positional encoding is not the same as context: A study on positional encoding for Sequential recommendation

### v1

@@
Abstract
The expansion of streaming media and e-commerce has led to a boom in recommendation systems, including Sequential recommendation systems, which consider the user’s previous interactions with items. In recent years, research has focused on architectural improvements such as transformer blocks and feature extraction that can augment model information. Among these features are context and attributes. Of particular importance is the temporal footprint, which is often considered part of the context and seen in previous publications as interchangeable with positional information. Other publications use positional encodings with little attention to them. In this paper, we analyse positional encodings, showing that they provide relative information between items that are not inferable from the temporal footprint. Furthermore, we evaluate different encodings and how they affect metrics and stability using Amazon datasets. We added some new encodings to help with these problems along the way. We found that we can reach new state-of-the-art results by finding the correct positional encoding, but more importantly, certain encodings stabilise the training.
1 Introduction
Sequential recommendation systems (SRS) have been booming within referral systems in the last few years since most of these systems are based on previous user behaviour and are present in almost all e-commerce, social media, and streaming platforms. Their effectiveness has led to various definitions of sequential recommendations and to their division into branches.
In general, sequential recommendation systems Fang et al. (2020), refer to recommendation systems (RS) where the interaction between the user and items is recorded sequentially. SRS are also related to session-based or session-aware recommendations. The latter two terms can be viewed as the sub-types of sequential recommendation when the interactions log is restricted to those made in the current session.
In this work, we adhere to the notation used in Fang et al. (2020) and use the term object to summarise the interactions. "Behavior object" means what the user interacts with, identified by an ID or a group of items. It may include text, images, and time. We will use "item(s)" for convenience. The interaction between items and users can use many different types of features. A typical split of those is attributes and context, where the attributes are features provided by the item, like a text description or an image, and the context refers to those features related to the interaction itself, like the time that a user tends to buy or search some streaming content.
A division based on the target task exists between the next-item recommendation and the next-basket recommendation Fang et al. (2020). In the next-item recommendation, a behaviour contains only one object (i.e., item), which could be a product, a song, a movie, or a location. In contrast, a behavior contains more than one object in next-basket recommendation.
When dealing with sequential systems, it is essential to consider the sequential nature of the data. Therefore, initial deep learning techniques utilised Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN) to address this issue. However, attentional models have become increasingly popular due to their effectiveness, leading to multiple approaches utilising this technique Wang et al. (2021). Currently, most state-of-the-art models use attentional networks.
@@ 1 Introduction
    We analyse existing work and highlight that more than the standard usage of $5$ runs is needed to ensure the reliability of the results.
  2. 2.
    We propose new encodings. Based on our first experiments with existing encodings, we propose new encodings to improve metrics and stability.
  3. 3.
@@
    We perform a detailed analysis of the types of encodings, their features (in-head, affected heads, etc.) and their ablations to see which factors improve the stability and the metrics.
  4. 4.
    We found correlations between the sparsity of the dataset and the encodings to get the best results. This precise data dependence, together with our study, provides a road map for the usage of encodings for SRS with transformer architectures.
2 Related Work
From an architectural perspective, SRS has evolved over the last two decades. Traditional methods in the field applied classic ML techniques, such as KNN for item-based systems (Davidson et al. (2010), Linden et al. (2003)), or applications of Markov Chains for different orders (He et al. (2017), He et al. (2016b), He and McAuley (2016a)). A common approach due to the sparsity of the interactions in RS is those based on matrix factorisation like DeepFM Guo et al. (2017) or CFM Xin et al. (2019). Among the different DL architectures, RNNs can capture sequence information and represent items based on their position in the list. One of the earliest and best known was Cho et al. (2014). The first approach from DL came using RNN, the most common architecture then. We can find models such as GRU4Rec Hidasi et al. (2016) or DREAM Yu et al. (2016). Another of the first approaches via DL was based on CNN models that can capture local features and consider time information in the input layer (Tang and Wang (2018), Tuan and Phuong (2017)). Among these were Caesar Li et al. (2021) and NextItNet Yuan et al. (2018). One of the first approaches using attention models was NARM Li et al. (2017), an encoder-decoder framework that uses attention at the bottleneck. The introduction of transformers and self-attention in sequential models came from AttRec Zhang et al. (2018).
Over time, different types of information have been identified to predict what a user might buy next. There are three main types of models for this: attribute-aware Rashed et al. (2022), context-aware Zhou et al. (2020), and time-aware Li et al. (2020). Attribute-aware models focus on the information provided by the items themselves. Context-aware models take into account the characteristics of the user’s interaction. Time-aware models are a subgroup of context-aware models that use the timing of the user’s interactions.
Using transformers in architecture has led to the introduction of positional embeddings. By default, the attention heads work like bags of tokens, and the positional information needs to be added to understand the sequential order of the tokens. Some approaches use time-aware information to replace the position, like in the case of CARCA Rashed et al. (2022). While this approach achieved excellent results, it faces some limitations. Replacing the positional embeddings with time stamps assumes that the order of the items will be implicitly deduced from this stamp. This can be more challenging to the model than simply an explicit representation. Furthermore, recent research in NLP Chen et al. (2021) has shown that introducing positional information within the attention, instead of the first head input, produces better results, which can not be done through the time stamps.
One of the first models considering the sequential information was GRU4Rec Hidasi et al. (2016), which concentrates on extracting information from the sequence of items instead of the user being an anonymous user model. A successful model from the context-aware point of view was DeepFM Guo et al. (2017), which extracted through the combination of factorisation models and deep learning extracts information from the context. Like GRU4Rec, using a self-attention layer to balance short-term intent and long-term preference can be found in SASRec Kang and McAuley (2018a). The extraction through a sequence of transformer encoders produces an output, which is compared with the target items through dot-product. This dot-product approach combines information from past behaviour and target representation, which was improved by an encoder-decoder architecture, where the encoder extracts past information, and the decoder combines the embedding representation for the target. This produced one of the SOTA models, CARCA Rashed et al. (2022). Another path of improvement from SASRec came from training the bidirectional transformer to infer the item-item relationship from the user’s historical interactions using the Cloze task in BERT4Rec Sun et al. (2019). Another one was TiSASRec Li et al. (2020), which further improves SASRec by considering time intervals between items.
We will focus on the position encoding applied to the most common datasets among these papers. For a deeper review of the different encoding techniques, one can see Zhao et al. (2024). Since there is no systematic study, each work addresses the encoding differently. In AttRec Zhang et al. (2018), they "furnish the query and key with time information by positional embeddings" using absolute embeddings. SASRec used absolute learnable position embeddings in the input sequence. They also tried the fixed position embedding as used in Vaswani et al. (2017) but found that this led to worse performance. In CARCA, which will be our base model, they tried using absolute encodings again, giving even worse results than no positional encoding. This disparity of results merits a more detailed analysis as in the present work.
3 Positional encodings
@@
Rotatory No APE Unl. Vector Encoding No Multi.
Rotatory + Con. Yes APE Unl. Vector Encoding Yes Joint
@@
Table 1: Encoding types: ’Type’ refers to absolute and relative. ’Learnable’ indicates whether the encoding parameters are learned during training. We also consider learnable to have a concatenation since there are learnable parameters. ’Range’ specifies the positions the encoding covers: ’Fix.’ fixed by the given length, ’Unl.’ means unlimited, while $4$ refers to the relative range. ’Layers’ represents the number of layers the encoding uses: "Vector Encoding" means they are added at the start by vectors (left in Fig: 1), All means in all the transformer blocks and First, at the first encoding block. ’Concatenation’ indicates whether there is a concatenation. ’Integration’ describes how the encoding is integrated with the model architecture: ’Add’ indicates a vector sum, ’Joint’ the concatenation, and ’Multi.’ for multiplication of matrices for those in the heads or element-wise for the others. Finally, None refers to no encoding.
@@
Figure 1: APE vs RPE : While APE encodings introduce the positional information as vectors before the $V$, $Q$ and $K$, projections, RPE add this information at the coefficient level in $K$ and $Q$. Parts affected by the position information appear in red.
3.1 Positional encodings
This section introduces the different types of encoding techniques used in this work. For a deeper review of the various encoding methods, check Zhao et al. (2024). Following Zhao et al. (2024), there are two prominent families of encoding methods. First, we have absolute positional encoding (APE ). These encodings are added before the first Transformer block, and the positional information provided is absolute, from the initial token to the last one. Second, there are the relative encoding methods (RPE ), which provide information about those items around the target item. These encodings are implemented inside the transformer’s head, more precisely in the Queries and Keys. In contrast to the common usage of encoding in Transformers for NLP tasks in RS, the encoding is reversed; the first item in the encoding order is the last one, and padding is applied on the left side Kang and McAuley (2018a). We chose this approach because the most relevant items are usually the ones that were recently visited or clicked. Due to the sinusoidal functions used in APE , the most pertinent items might be at the start because the assigned frequency becomes less distinguishable the later an item appears. We did some preliminary experiments to check if this is also right for another encoding, and the reversed encoding always produced better results, so in what follows, all the encoding will be considered as reversed.

### v2

@@
Abstract.
The rapid growth of streaming media and e-commerce has driven advancements in recommendation systems, particularly Sequential Recommendation Systems (SRS). These systems employ users’ interaction histories to predict future preferences. While recent research has focused on architectural innovations like transformer blocks and feature extraction, positional encodings, crucial for capturing temporal patterns, have received less attention. These encodings are often conflated with contextual, such as the temporal footprint, which previous works tend to treat as interchangeable with positional information. This paper highlights the critical distinction between temporal footprint and positional encodings, demonstrating that the latter offers unique relational cues between items, which the temporal footprint alone cannot provide. Through extensive experimentation on eight Amazon datasets and subsets, we assess the impact of various encodings on performance metrics and training stability. We introduce new positional encodings and investigate integration strategies that improve both metrics and stability, surpassing state-of-the-art results at the time of this work’s initial preprint. Importantly, we demonstrate that selecting the appropriate encoding is not only key to better performance but also essential for building robust, reliable SRS models.
copyright: noneconference: ; ; price: doi: isbn:
1. Introduction
Sequential Recommendation Systems (SRS) have become integral to personalizing user experiences by predicting the next item of interest based on historical interaction sequences. These systems are widely deployed across e-commerce, social media, and streaming platforms and significantly enhance user engagement and satisfaction. As defined in (Fang et al., 2020), SRS records user-item interactions sequentially, closely related to session-based and session-aware recommendations, which focus specifically on interactions within a single session.
Figure 1. Illustration of our proposed Rotatory  embeddings. Similar to RoPE , it encodes positional information through the angles of the embeddings. However, unlike RoPE , which assigns angles proportionally based on position, Rotatory learns these angles during training. It allows the model to determine the most relevant positional information and adjust their significance accordingly.
Capturing the order of data is essential in sequential systems. Initially, Deep Learning techniques utilized Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN) to model sequential dependencies. However, attentional-based models have gained prominence due to their superior effectiveness, leading to numerous self-attention approaches (Wang et al., 2021). Most state-of-the-art (SOTA) models in SRS now leverage attentional-based networks. A critical component of these systems is positional encoding (PE), which imparts information about the order of items within a sequence. The choice and implementation of PEs are crucial for optimizing performance and stability.
Despite the widespread adoption of various PE techniques, there is a noticeable gap in research concerning the stability of model performance across different runs. However, many existing publications focus on enhancing performance metrics without adequately addressing the underlying stability issues introduced by different encoding strategies. Other studies underestimate the relevance of positional encoding, considering them part of the context.
To bridge this gap, we aim to investigate the stability problem in SRS models through positional encodings. We conduct extensive experiments to evaluate how different positional encodings influence these models’ performance and stability. Building on this foundation, we introduce a novel positional encoding method termed Rotatory , along with its concatenated variant, Rotatory + Con , and ablations of the previous encoding method.
@@ 1. Introduction
    We analyze existing work and highlight that more than the standard usage of $5$ runs is needed to ensure the reliability of results.
  2. (2)
@@
    We conduct a detailed analysis of various encoding types, their features (such as in-head and affected heads), and perform ablation studies to identify factors that improve stability and performance.
  3. (3)
    We introduce a novel rotatory-based encoding method, alongside integration techniques, such as concatenation, which have not been explored previously. These contributions aim to improve both performance metrics and training stability.
  4. (4)
    Our findings show that encoding effectiveness is primarily influenced by dataset deviation, rather than sparsity or data type. This insight, combined with our analysis, offers a clearer guide for using encodings in SRS with transformer architectures.
2. Related Work
@@
Rotatory Yes APE Unl. Vector Encoding No Add
Rotatory + Con Yes APE Unl. Vector Encoding Yes Joint
@@
Table 1. Encoding types: ’Type’ specifies absolute or relative. ’Learnable’ indicates if encoding parameters are trained, including concatenation for learnable encodings. ’Range’ denotes position coverage: ’Fix.’ for fixed length, ’Unl.’ for unlimited, and $4$ for relative range. ’Layers’ refers to encoding application: ’Vector Encoding’ adds vectors at the start (left in Fig 3), ’All’ applies to all transformer blocks, and ’First’ to the first block. ’Concatenation’ shows if concatenation is used. ’Integration’ describes the method: ’Add’ for vector sum, ’Joint’ for concatenation, and ’Multi.’ for matrix multiplication in heads or element-wise otherwise. ’None ’ means no encoding.
2.1. Architectures & Encoding in SRS
From an architectural perspective, SRS have significantly evolved over the past two decades. Early approaches employed K-Nearest Neighbors (KNN) for item-based recommendations (Davidson et al., 2010; Linden et al., 2003) and Markov Chains for modeling sequential interactions (He et al., 2017, 2016b; He and McAuley, 2016a). To address interaction sparsity, matrix factorization methods like DeepFM (Guo et al., 2017) and CFM (Xin et al., 2019) were introduced, enhancing the ability to handle sparse data effectively.
With the advent of Deep Learning, RNNs became popular for capturing sequence information and item positions, exemplified by models such as GRU4Rec (Hidasi et al., 2016) and DREAM (Yu et al., 2016). Concurrently, CNNs were utilized to capture local features and temporal information, as seen in Caesar (Li et al., 2021) and NextItNet (Yuan et al., 2018), among others (Tang and Wang, 2018; Tuan and Phuong, 2017). Attention mechanisms were first introduced in NARM (Li et al., 2017), and transformers with self-attention were pioneered in SRS by AttRec (Zhang et al., 2018).
Different types of models are used to predict user behavior: attribute-aware (Rashed et al., 2022), context-aware (Zhou et al., 2020), and time-aware (Li et al., 2020). These models focus on item information, user interaction characteristics, and the timing of interactions, respectively. The integration of transformers has introduced positional embeddings. Some approaches, like CARCA (Rashed et al., 2022), use time-aware information to replace traditional positional encodings, assuming that the order of items can be deduced from timestamps. However, recent NLP research (Chen et al., 2021) has shown that incorporating positional information directly within the attention mechanism, rather than relying on timestamps, yields better results.
One of the initial models to capture sequential information was GRU4Rec (Hidasi et al., 2016), which extracts data from item sequences. The context-aware DeepFM (Guo et al., 2017) combines factorization methods with deep learning to leverage contextual information. Similarly, SASRec (Kang and McAuley, 2018a) employs a self-attention layer to balance short-term intent and long-term preferences, using transformer encoders to compare past behavior with target items via dot-products. This encoder-decoder architecture, which extracts past information while synthesizing target embeddings, led to the development of CARCA (Rashed et al., 2022).
We focus on positional encoding across various datasets used in these studies. Each work approaches encoding differently: AttRec (Zhang et al., 2018) employs absolute embeddings to incorporate time information into the query and key, while SASRec utilizes absolute learnable position embeddings in the input sequence. CARCA experimented with absolute encodings but observed worse performance compared to using no positional encoding. Recent advancements, such as EulerFormer (Tian et al., 2024), introduce complex vector attention to integrate semantic and positional information. This variety in encoding strategies highlights the need for a systematic analysis, which we provide in this work.
Figure 2. (1) Vector encoding added before Transformer blocks, (2) First head RPE: Relative encoding applied only to the first block (RopeOne ), and (3) All head RPE: Relative encoding integrated into every block (RMHA-4 ).
@@ 2.1. Architectures & Encoding in SRS
Figure 3. APE vs RPE : While APE encodings introduce the positional information as vectors before the $V$, $Q$ and $K$. RPE add this information at the coefficient level to $K$ and $Q$. Parts affected by the position information appear in red.
2.2. Positional encodings
Positional encoding enables transformer-based models to incorporate item order within sequences. Encoding methods can be categorized into absolute and relative encodings (Zhao et al., 2024) (Figure 3). In Table 1, we have a detailed classification of the main encodings in this paper.
Absolute Positional Encoding (APE )
They assign a unique position vector to each item based on its position in the sequence. These encodings are added before the first Transformer block, providing absolute positional information. Models like SASRec (Kang and McAuley, 2018a) utilize learnable absolute position embeddings, allowing the model to learn position representations during training. There are two common variations of absolute positional encodings: fixed (Abs ) and learnable (Learnt ).
Absolute Positional Encoding (Abs )
Absolute positional encoding adds fixed values to item embeddings based on their absolute position in the sequence. The most common approach uses sinusoidal functions, as seen in BERT (Devlin et al., 2019). This encoding is added at the input layer as part of the item encoding (Figure, left 3).
The encoding for an item at position $pos$ in a sequence is calculated using sine and cosine functions:
(1) $PE_{(pos,2i)}=\sin\left(\frac{{pos}}{{10000^{(2i/d)}}}\right)$
(2) $PE_{(pos,2i+1)}=\cos\left(\frac{{pos}}{{10000^{(2i/d)}}}\right)$
where $i$ is the dimension, and $d$ is the embedding dimension. This encoding is added to the original embedding:
(3) $x=x+PE_{(pos,*)}$
Learnable Encoding (Learnt )
Learnable position embeddings treat positional information as trainable parameters. Unlike fixed encodings, the model learns optimal position representations based on the data context, denoted as learnable encoding (Learnt ). Positional encodings are added to the input sequence embeddings.
Relative Positional Encoding (RPE )
Relative Positional Encoding (Figure 3, right) encodes the relative distances between items instead of their absolute positions, enabling models to better capture contextual relationships within user interactions. Unlike Absolute Positional Encoding (APE ), which requires fixed input lengths and provides positional information from the start to the end of the sequence, RPE offers flexibility by focusing on the relationships between items regardless of their absolute positions. A downside of these positional encoding mechanisms is that they are significantly slower during training and inference since they add an extra step in the self-attention layer (Press et al., 2022).